Session: Nov 1998 (session 4)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Zacchiroli at Sun Nov 22 20:20:33 MET 1998) Success
No known problems, all the scans recorded.

Noto (C.Trigilio at Mon Nov 23 10:20:44 MET 1998) Success

Torun (K.Borkowski at Mon Nov 23 10:35:20 MET 1998) Success
Between 17:00 and 17:40 UTC a few scans recorded off-source, because the source was too close to or below the horizon.

Effelsberg (dgraham at Mon Nov 23 11:50:11 MET 1998) Minor failures
Tsys often measured while slewing. Since Tcal is only 1.5K, these Tsys values are kaputt

Westerbork (Tony Foley at Tue Nov 24 16:29:48 MET 1998) Minor failures
Basically looks OK, but one of the 8 cryogenic frontends had warmed up.

Jodrell (Alastair Gunn at Wed Nov 25 07:47:57 MET 1998) Minor failures
Problems with tape and terminal at start of schedule. Schedule running fine at 12:49 UT.

Cambridge (Alastair Gunn at Wed Nov 25 08:18:35 MET 1998) Success
No known problems.

Onsala (A. Polatidis at Wed Nov 25 16:20:51 MET 1998) Success
Strong winds but we could observe. No problems with the data.

Calibration (Jean\-Francois Desmurs at Fri Dec 11 16:16:55 MET 1998) Minor failures
(Bad values for JB and no value for CM). The ANTAB formatted file for this experiment can be found on VLBEER in the nov98 directory under the name: eg019.antab. Read the comments in the file to get further information.